OpenStudy (anonymous):

How do I solve for y 5x-3y=-3 2x-6y=-6

OpenStudy (saifoo.khan):

multiply the first equation by 2

OpenStudy (anonymous):

-1 1 0 2

OpenStudy (anonymous):

y=-1x+1?

OpenStudy (saifoo.khan):

1

OpenStudy (anonymous):

can you show me how?

OpenStudy (saifoo.khan):

my calc sucks at showing the workings! :/ let me do it for u.

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

i just want to know for future reference.

OpenStudy (saifoo.khan):

10x-6y=-6 2x-6y=-6 -10x +6y = 6 2x -6y = -6 ======== -8x = 0 x = 0

OpenStudy (saifoo.khan):

Now, x = 0 in any equation will give u 1

OpenStudy (saifoo.khan):

And sorry, multiply the equation by -2.

OpenStudy (anonymous):

Thanks... when i multiplied the equation i didnt get it... thanks :)

OpenStudy (saifoo.khan):

Oh yea there was a mistake!!